OpenStudy (legomyego180):

Use differentials to find the approximate value:

OpenStudy (legomyego180):

\[203^{\frac{ 1 }{ 2 }} 18^{\frac{1}{4}}\]

OpenStudy (legomyego180):

Or, a different question: Minimize: \[x^2+y^2+z^2=7\]

OpenStudy (solomonzelman):

Example: Approximate \(\sqrt{126}\) using differentials. ------------------------------------------------- \(\sqrt{126}\) is near \(\sqrt{121}=11\), so we will set \(f(x)=\sqrt{x}\) with \(x=121\) and \(\Delta x=5\). \(f(x+\Delta x)=f(x)+\Delta y~{\color{red}{=}}~f(x)+f'(x)\Delta x\) \(f(x+\Delta x)=f(x)+(2\sqrt{x})^{-1}\Delta x\) \(\sqrt{126}=\sqrt{121}+\frac{1}{2\sqrt{121}}(5)=11+\frac{5}{22}\)

OpenStudy (solomonzelman):

The red equal sign is the sign that corresponds to where I made my approximation. Precisely, we are approximating by assuming \(\Delta y \approx f'(x) \Delta x\). (The concept is basically the same as with the tangent line approximation.)

OpenStudy (solomonzelman):

So, in your case, you can approximate \(\sqrt{203}\), by setting \(\tiny \\[0.5em]\) (1) \(f(x)=\displaystyle\sqrt{x}\), \(\tiny \\[0.5em]\) (2) \(x=196\) (because \(\sqrt{196}=14\)), and \(\tiny \\[0.5em]\) (3) \(\Delta x=7\) (because \(203-196=7\))

OpenStudy (solomonzelman):

Then, you can approximate \(\sqrt[4]{18}\), by setting \(\tiny \\[0.5em]\) (1) \(f(x)=\displaystyle\sqrt[4]{x}\), \(\tiny \\[0.5em]\) (2) \(x=16\) (because \(\sqrt[4]{16}=2\)), and \(\tiny \\[0.5em]\) (3) \(\Delta x=2\) (because \(18-16=2\)) Note that \(\Delta y=f'(x)\Delta x=\displaystyle \frac{\Delta x}{4\sqrt[4]{x^3}}= \frac{2}{4\sqrt[4]{16^3}}= \frac{1}{2\sqrt[4]{(2^4)^3}}= \frac{1}{16}\).

OpenStudy (solomonzelman):

Then, as you approximate \(\sqrt{203}\) and \(\sqrt[4]{18}\) individually multiply these individual approximations to get the approximation for the product.

OpenStudy (solomonzelman):

Next, I will minimize \(z^2+x^2+y^2=13\). The \(z\)-value is the smallest when \(z=-\sqrt{13}\), because \(\sqrt{13}\) is the smallest value on the sphere. So we already know what the minimum is. So we need \(x^2+y^2=0\). The only way for \(x\) and \(y\) to meet this equation is \((x,y)=(0,0)\). So the absolute minimum is \(x^2+y^2+z^2=13\). (No calculus here.)

OpenStudy (solomonzelman):

The absolute minimum occurs \((0,0,-\sqrt{13})\) .... For any sphere with radius \(\delta\), the minimum is \((0,0,-\delta)\), and the maximum is \((0,0,\delta)\).

OpenStudy (solomonzelman):

well, for any such sphere that is centered at the origin of course.

OpenStudy (legomyego180):

Im here btw, just working this out alongside you

OpenStudy (solomonzelman):

I am pretty sure you understand the sphere thing. You know what the sphere looks like precisely, so there you know the maximum and the minimum of it ....

OpenStudy (solomonzelman):

Do you have any questions about the approximation using differentials?

OpenStudy (legomyego180):

\(\color{blue}{\text{Originally Posted by}}\) @SolomonZelman Example: Approximate \(\sqrt{126}\) using differentials. ------------------------------------------------- \(\sqrt{126}\) is near \(\sqrt{121}=11\), so we will set \(f(x)=\sqrt{x}\) with \(x=121\) and \(\Delta x=5\). \(f(x+\Delta x)=f(x)+\Delta y~{\color{red}{=}}~f(x)+f'(x)\Delta x\) \(f(x+\Delta x)=f(x)+(2\sqrt{x})^{-1}\Delta x\) \(\sqrt{126}=\sqrt{121}+\frac{1}{2\sqrt{121}}(5)=11+\frac{5}{22}\) \(\color{blue}{\text{End of Quote}}\) Is the delta y here supposed to be delta x?

OpenStudy (solomonzelman):

Oh, no it shouldn't.

OpenStudy (legomyego180):

that was the only thing that threw me off.

OpenStudy (solomonzelman):